# Localizing Before Answering: A Hallucination Evaluation Benchmark for Grounded Medical Multimodal LLMs

### v2

@@ 1 Introduction
Large Multimodal Models (LMMs), or Multi-modal Large Language Models (M-LLMs) Alayrac et al. (2022); Bai et al. (2023); Liu et al. (2024b) such as GPT-4V OpenAI et al. (2023) achieve superb performance in understanding data from multiple modalities, such as vision, and language, and generating human-like texts. Leveraging the strong capabilities of those foundational models, recent works Li et al. (2024); Chen et al. (2024); Zhang et al. (2023a); Chen et al. (2023); Moor et al. (2023); Wu et al. (2023b, a); Yan and Pei (2022) have developed M-LLMs for medical imaging applications such as medical Visual Question Answering (VQA). Recently, many multimodal foundation models (LLaVA-Med Li et al. (2024), CheXagent Chen et al. (2024), GPT-4 OpenAI et al. (2023), Gemini Team et al. (2023)) have emerged and demonstrated impressive reasoning and comprehension capabilities on biomedical queries and related domains. These models can assist clinical professionals in interpreting and gaining insights on many medical images, helping the diagnosis and treatment processes be more efficient.
@@
To address these concerns, we introduce a new benchmark, dubbed HEAL-MedVQA (Hallucination EvAluation via Localization in Medical VQA) to dissect the LLM’s ability to localize at visual evidence when answering. Specifically, this paper proposes two evaluation protocols, called Textual Perturbation Test (TPT), and Visual Perturbation Test (VPT), which probe LLMs’ sensitivity towards textual and visual shortcuts. To support this evaluation, we curate adversarial VQAs and procure pixel-level annotations of affected anatomy regions from three board-certified radiologists from two large-scale public datasets, MIMIC-CXR Johnson et al. (2019) and VinDr-CXR Nguyen et al. (2020). In total, we construct over 60,000 question-answer pairs probing LLM’s textual and visual shortcut learning. Our curated medical benchmark provides structured codebase, facilitating comprehensive comparisons of advanced LLMs.
@@ 3.1 New Evaluation Protocol — Sensitivity Test
  • Vision Shortcut: The model might concentrate on non-relevant regions (as illustrated in Fig. 1a). This behavior results in the unreliability of LMMs, even when providing correct answers, as they fail to rely on medical visual evidence.
@@
Figure 2: Textual and Visual Perturbation Tests for our experiments. For Textual Perturbation Test (TPT), we randomly replace either the anatomy or the disease in the original question with another anatomy/disease so that the new disease pair is logical (present in the training dataset). For Visual Perturbation Test (VPT), we replace the visual area of interest (e.g. left lower lung) with the segmented area of the same region from another image that does not share the same diagnosis as the current one. We measure the percentage of instances a model changes the response from “Yes” to “No” to evaluate the model’s sensitivity to visual and language shortcuts.
@@
  • We randomly replace either the anatomy or disease with another new anatomy/or disease. The new anatomy and disease pair has to be logical, which means they does co-occur in the dataset. For example, as shown in Fig. 2 (Above), the question “Does left lower lung suffer from pneumonia?” can be replaced by either switching the anatomy to right lower lung, or the disease to pulmonary fibrosis.
@@ 3.2 Dataset Curation
To assess the LLM’s robustness against both textual and visual shortcuts, we introduce a new benchmark called HEAL-MedVQA (Hallucination Evaluation via Localization in Medical VQA). Built upon the two large scale radiology datasets, VinDr-CXR Nguyen et al. (2020) and MIMIC-CXR Johnson et al. (2019), we construct HEAL-MedVQA by mapping spatial relationships between diseases and anatomical locations, and probing the model’s robustness towards disease and location perturbation on both texts and images. Fig. 4 illustrates the overall pipeline of our data construction. We aim to extract the disease bounding boxes, and anatomical masks from the image. Based on the overlapping between disease and anatomy localization, we determine their spatial relationships. Then several QA templates are used to generate VQA samples.
@@
To get the anatomy-disease relation, we measure the overlapping between the disease bounding box and anatomical segmentation mask. If the IoU score between a disease-anatomy pair is over a threshold $\delta=0.5$, then that disease occurs on the mentioned anatomy.
@@
Table 1: Comparison of Med-VQA datasets. indicates that visual annotations are for all anatomic regions in the image instead of the one in the question.
@@ 4 Proposed Method
Figure 5: The proposed LoBA frameworks consists of two phases. Grounded text generation: during training, our model learns to localize and segment a region of interest mentioned in the question. Self-prompting: during inferencing, the model attention is calibrated to attend to the segmented region. With the highlighted attention, the model refines and generates visually grounded answers that are more robust to hallucinations.
@@ 4.1 Grounded Text Generation
We use a <SEG> token, denoting the token that will be used in later steps to obtain the segmentation of the relevant anatomies. Using this token, we instruct LLMs to localize on visual evidence before answering. As such, we format the response as “The location of {relevant areas} is at <SEG>”. For example, if the question is “Are there any abnormalities at the right lower lung?”, the output of the LMM will be “The location of the right lower lung is at <SEG>”. Note that, in our format, the model has to produce the visual evidence first, before giving the final answers.
The LLM is trained to generate the above response:
$\hat{y}_{text}=LLM(x_{text},x_{img}).$ (1)
We use the projected hidden state $h_{seg}$ from the <SEG> token as the query to a the segmentation model. This can be formulated as $\hat{M}=Seg(x_{img},h_{seg})$. Our LLM is trained via the text objectives. Similar to other LLMs Touvron et al. (2023); OpenAI et al. (2023), we formulate the text generation loss as the cross entropy loss between the generated text and the ground truth
$\mathcal{L}_{text}=CE(\hat{y}_{text},y_{text}).$ (2)
@@
$\mathcal{L}_{seg}=\lambda_{BCE}BCE(\hat{M},M)+\lambda_{DICE}DICE(\hat{M},M).$ (3)
@@
$\mathcal{L}=\lambda_{text}\mathcal{L}_{text}+\lambda_{seg}\mathcal{L}_{seg}.$ (4)
@@ 4.2 Attention Highlighter via Self-Prompting
While the proposed training paradigm allows the model to localize the region of interests when answering, hallucination can still occur when the model pays little attention to the image compared to the text. To deal with this, we propose to highlight the attention on the queried regions by self-prompting the segmentation map. Our module consists of two functions: attention reweighting and contrastive decoding.
@@
Model MIMIC VinDr
Yes/No Open-ended Yes/No Open-ended
F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall
Proprietary Models
Gemini-1.5-Flash-8B Team et al. (2023) 0.548 0.520 0.569 0.091 0.127 0.024 0.571 0.604 0.543 0.038 0.098 0.023
GPT-4-O OpenAI et al. (2023) 0.646 0.706 0.556 0.109 0.143 0.089 0.599 0.660 0.487 0.097 0.087 0.069
GPT-4-Vision OpenAI et al. (2023) 0.521 0.652 0.409 0.087 0.093 0.067 0.568 0.637 0.424 0.074 0.053 0.072
Open-source Models
CheXagent Chen et al. (2024) 0.677 0.621 0.695 0.425 0.450 0.416 0.610 0.628 0.603 0.374 0.342 0.330
BioMedGPT Zhang et al. (2023a) 0.708 0.734 0.672 0.397 0.387 0.412 0.624 0.606 0.631 0.289 0.265 0.310
MedFlamingo Alayrac et al. (2022) 0.562 0.574 0.552 0.365 0.390 0.310 0.584 0.600 0.547 0.301 0.285 0.322
LLaVA-Med Li et al. (2024) 0.722 0.712 0.709 0.551 0.541 0.565 0.703 0.708 0.699 0.523 0.519 0.530
LISA Lai et al. (2024) 0.716 0.692 0.728 0.524 0.556 0.511 0.707 0.712 0.694 0.521 0.509 0.534
Adversarial Techniques
PH + LLaVA-Med Zhang et al. (2024b) 0.735 0.754 0.724 0.563 0.589 0.540 0.716 0.725 0.698 0.522 0.510 0.527
VCD + LLaVA-Med Leng et al. (2024) 0.716 0.708 0.721 0.566 0.583 0.538 0.652 0.699 0.591 0.501 0.547 0.497
CRG + LLaVA-Med Zhang et al. (2022) 0.731 0.715 0.725 0.544 0.563 0.554 0.689 0.667 0.702 0.509 0.524 0.494
LobA w/o self-prompt 0.727 0.716 0.731 0.544 0.557 0.535 0.724 0.745 0.706 0.511 0.526 0.507
LobA (Ours) 0.752 0.759 0.743 0.581 0.593 0.572 0.728 0.758 0.711 0.542 0.533 0.560
Table 2: Performance comparison across different models on the proposed HEAL-MedVQA benchmark on MIMIC and VinDr datasets for Yes/No and Open-ended VQA tasks. Text in bold and underline highlights the best and second best results, respectively.
@@ 5 Experimental Results
[…]
Textual Perturbation Test. Table 3 reports the sensitivity of each model to textual language bias, measuring the percentage of changed True Positive answers when either switching an anatomy or a disease. The higher the percentage is, the more robust a model is from textual shortcut bias. Our proposed framework achieves the highest sensitivity measure overall, with the ratio of changed answers when changing anatomy/disease to be around 0.77 to 0.80. Enhancing the attention to the queried regions as in our proposal helps alleviate the textual and visual shortcut learning. Lastly, fine-tuned open-source models are generally more robust than closed-source models like GPT-4-O, showing that fine-tuning LLMs on our adversarial data remedies their hallucination problems.
[…]
Ablation studies. Table 5 reports the effects of each component in our LobA framework: the LLaVA Med backbone, Grounded Text Generation module, and the full framework with self-prompting. Applying Grounded Text Generation improves the model’s performance in some categories, showing the effectiveness of localization-aware training.With the self-prompting framework, the accuracy and textual/visual robustness improves greatly, most evidently from 0.763 to 0.797 in TPT score, or 0.722 to 0.752 in MIMIC-CXR dataset Yes/No F1 Score.
Qualitative analysis. Fig 11 showcases our LobA and some other LMM’s response on a few qualitative examples, as well as the visual attention map on the input image. We can see that LobA manages to query the attention weights much more intuitively compared to LLaVA-Med or LISA While LLaVA-Med and LISA’s attention are scattered into many irrelevant areas, in the first example LobA’s main attention weights are allocated around the right lower lung area, and the second example is near the left upper lung. Thanks to the grounding generation module, followed by LobA’s self prompting module to automatically re-weight the attention to the area of interest, the model is able to have better visual reasoning capabilities on the area of interest, leading to more accurate answers compared with other methods.
@@ 6 Conclusion
Appendix A Computational Efficiency
We present the inference time (seconds) and GPU consumption (MB) of LoBA on a system with a single A100 80GB GPU for 100 samples in comparison with some other medical LMMs in Tab. 6. Our LobA yields computational efficiency comparable to that of existing methods, while achieving superior localization capabilities.
CheXagent Med-Flamingo LLaVA-Med LobA (Ours)
Inf. Time 113.4 126.5 87.0 101.2
GPU Mem 23406 26453 17729 25254
Table 6: Comparison of inference time and GPU consumption
Appendix B Benchmarking Other General LLMs
We report the F1 score of additional proprietary and open-source LMMs (fine-tuned) on the VinDr subset in Tab. 7
Model Gemini-1.5 Claude 3.7 Qwen 2.5 VL InternVL LobA
Pro Sonnet 7B-Instruct 2.5 8B (Ours)
Yes/No 0.583 0.610 0.660 0.654 0.728
Open 0.078 0.062 0.413 0.378 0.542
Table 7: Performance on VinDr for general LMMs
Appendix C Ablation study
We conducted two ablation studies, examining the effect of added attention weights $\beta$ (Tab. 8) and the effect of segmentation quality on the model’s results (Fig 7).
In summary, low weights cause the segmented area to have minimal influence, harming results, while excessive weight (e.g., 4.0) shifts focus away from the question, also degrading performance. For the second ablation study, the quality of the segmentation indeed correlates directly with LoBA’s performance. Overall, our model outperforms LISA baselines when the segmentation Dice score $>0.25$, which accounts for more than $10\%$ of the test set. This shows that our design is already beneficial even when the segmentation module can localize the rough regions to attend to. For samples with Dice $<0.25$, which only account for 9.47% of data, our LoBA yields a slight performance decrease.
$\beta$ Yes/No F1 Open F1
0.5 0.694 0.476
1 0.698 0.510
2 0.720 0.545
3 0.728 0.542
4 0.712 0.501
Table 8: Effect of $\beta$ on performance on VinDr
Figure 7: Yes/No F1 w.r.t segmentation qualities. Data portion (%) on test set is also shown.
Appendix D LobA Prompt Template
Figure 8: LobA Instruction Template
Figure 9: GPT and Gemini Models Instruction Template for Yes/No questions
Figure 10: GPT and Gemini Models Instruction Template for Open-ended questions
Figure 11: Qualitative Examples from our dataset
We propose a Grounded Text Generation training paradigm to enable LMM to localize the area of interest with the segmentation token “[SEG]“, before giving its answer for the model to be aware of the localization information.
As such, our model is fine-tuned to follow a strict template as “The location of {relevant areas} is at <SEG>. The answer is …”. To make sure the model’s output comply with this standard, we designed a few-shot instruction template for LobA so that the model can follow the desired template more accurately, displayed in Fig 8
Appendix E Proprietary Models Template
For GPT-4-O and GPT-4-Vision and Gemini-1.5-Flash-8B models, we modify the template of these proprietary models for yes/no questions and open-ended questions separately to allow them to follow our instructions correctly. These 2 templates are displayed by Fig 10 and 9
Appendix F Evaluation Template
Figure 12: Yes/No evaluation prompt template
Figure 13: Open Ended evaluation prompt template
We utilize LLaMA-3 as the LLM backbone to extract the labels from the model’s natural language answers with the following prompt templates, Fig 12 and Fig 13
Appendix G Question example and qualitative results
We present some qualitative examples, as well as some model’s response and their visual attention map in Fig 11.

### v3

@@ 1 Introduction
Large Multimodal Models (LMMs), or Multi-modal Large Language Models (M-LLMs) Alayrac et al. (2022); Bai et al. (2023); Liu et al. (2024b) such as GPT-4V OpenAI et al. (2023) achieve superb performance in understanding data from multiple modalities, such as vision, and language, and generating human-like texts. Leveraging the strong capabilities of those foundational models, recent works Li et al. (2024); Chen et al. (2024); Zhang et al. (2023a); Chen et al. (2023); Moor et al. (2023); Wu et al. (2023b, a); Yan and Pei (2022) have developed M-LLMs for medical imaging applications such as medical Visual Question Answering (VQA). Recently, many multimodal foundation models (LLaVA-Med Li et al. (2024), CheXagent Chen et al. (2024), GPT-4 OpenAI et al. (2023), Gemini Team et al. (2023)) have emerged and demonstrated impressive reasoning and comprehension capabilities on biomedical queries and related domains. These models can assist clinical professionals in interpreting and gaining insights on medical images, helping the diagnosis and treatment processes be more efficient.
@@
To address these concerns, we introduce a new benchmark, dubbed HEAL-MedVQA (Hallucination EvAluation via Localization in Medical VQA) to dissect the LLM’s ability to localize at visual evidence when answering. Specifically, this paper proposes two evaluation protocols, called Textual Perturbation Test (TPT), and Visual Perturbation Test (VPT), which probe LLMs’ sensitivity towards textual and visual shortcuts. To support this evaluation, we curate adversarial VQAs and procure pixel-level annotations of affected anatomy regions from three board-certified radiologists from two large-scale public datasets, MIMIC-CXR Johnson et al. (2019) and VinDr-CXR Nguyen et al. (2020). In total, we construct over 60,000 question-answer pairs probing LLM’s textual and visual shortcut learning. Our curated medical benchmark provides structured codebase, facilitating comprehensive comparisons of advanced LLMs. The code and dataset will be released at https://loba-ijcai25.github.io.
@@ 3.1 New Evaluation Protocol — Sensitivity Test
  • Vision Shortcut: The model might concentrate on non-relevant regions (as in Fig. 1a). This behavior results in the unreliability of LMMs, even when providing correct answers, as they fail to rely on medical visual evidence.
@@
Figure 2: Textual and Visual Perturbation Tests for our experiments. For Textual Perturbation Test (TPT), we randomly replace either the anatomy or the disease in the original question with another so that the new anatomy/disease pair is logical (present in the training dataset). For Visual Perturbation Test (VPT), we replace the visual area of interest (e.g. left lower lung) with the segmented area of the same region from another image that does not share the same diagnosis as the current one. We measure the percentage of instances a model changes the response from “Yes” to “No” to evaluate the model’s sensitivity to visual and language shortcuts.
@@
  • We randomly replace either the anatomy or disease with another new anatomy/or disease. The new anatomy and disease pair has to be logical, which means they does co-occur in the dataset. For example, in Fig. 2 (Above), the question “Does left lower lung suffer from pneumonia?” can be replaced by either switching the anatomy to right upper lung, or the disease to pulmonary fibrosis.
@@ 3.2 Dataset Curation
To assess the LLM’s robustness against both textual and visual shortcuts, we introduce a new benchmark called HEAL-MedVQA (Hallucination Evaluation via Localization in Medical VQA). Built upon the two large scale radiology datasets, VinDr-CXR Nguyen et al. (2020) and MIMIC-CXR Johnson et al. (2019), we construct HEAL-MedVQA by mapping spatial relationships between diseases and anatomical locations, and probing the model’s robustness towards disease and location perturbation on both texts and images. Fig. 4 illustrates our overall pipeline of data construction. We extract the disease bounding boxes and anatomical masks from the image, and based on the overlap between disease and anatomy localization, we determine their spatial relationships. Then several QA templates are used to generate VQA samples.
@@
[…]
@@
Table 1: Comparison of Med-VQA datasets. indicates visual annotations for all anatomic regions in the image instead of the one questioned.
@@ 4 Proposed Method
Figure 5: The proposed LoBA frameworks consists of two phases. Grounded text generation: during training, our model learns to localize and segment a region of interest mentioned in the question. Self-prompting: during inferencing, the model attention is calibrated to attend to the segmented region. With the attention, the model refines and generates visually grounded answers that are more robust to hallucinations.
@@ 4.1 Grounded Text Generation
We use a <SEG> token, denoting the token that will be used in later steps to obtain the segmentation of the relevant anatomies. Using this token, we instruct LLMs to localize on visual evidence before answering. As such, we format the response as “The location of {relevant areas} is at <SEG>”. For example, if the question is “Are there any abnormalities at the right lower lung?”, the output of the LMM will be “The location of the right lower lung is at <SEG>”. Note that, in our format, the model has to produce the visual evidence first, before giving the final answers. The LLM is trained to generate the above response:
$\hat{y}_{\text{text}}=LLM(x_{\text{text}},x_{\text{img}}).$ (1)
We use the projected hidden state $h_{\text{seg}}$ from the <SEG> token as the query to a the segmentation model. This can be formulated as $\hat{M}=Seg(x_{\text{img}},h_{\text{seg}})$. Our LLM is trained via the text objectives. Similar to other LLMs Touvron et al. (2023); OpenAI et al. (2023), we formulate the text generation loss as the cross entropy loss between the generated text and the ground truth
$\mathcal{L}_{text}=\text{CE}(\hat{y}_{\text{text}},y_{\text{text}}).$ (2)
@@
$\mathcal{L}_{seg}=\lambda_{\text{BCE}}\text{BCE}(\hat{M},M)+\lambda_{\text{DICE}}\text{DICE}(\hat{M},M).$ (3)
@@
$\mathcal{L}=\lambda_{\text{text}}\mathcal{L}_{\text{text}}+\lambda_{\text{seg}}\mathcal{L}_{\text{seg}}.$ (4)
@@ 4.2 Attention Highlighter via Self-Prompting
While the proposed training paradigm allows the model to localize the region of interest when answering, hallucination can still occur when the model pays little attention to the image compared to the text. To deal with this, we propose to highlight the attention on the queried regions by self-prompting the segmentation map. Our module consists of two functions: attention reweighting and contrastive decoding.
@@ 5 Experimental Results
Experimental Results and Discussion. VQA Accuracy. Table LABEL:tab:performance_comparison reports the quantitative accuracy of model’s answers on MIMIC and VinDr datasets. Proprietary models generally achieve sub-optimal accuracy, with the accuracy being less than 50% for most models on binary Yes/No questions and mostly less than 10% for open-ended questions. This shows the challenging nature of our HEAL-MedVQA benchmark, requiring the model to be robust against visual and textual shortcut learning. Our proposed framework consistently outperforms recent advances in M-LLMs and adversarial VQA techniques even after fine-tuning them. Notably, the proposed LoBA outperforms the state-of-the-art baseline Prompt Highlighter Zhang et al. (2024b) by 3.19% and 3.63% on Open-ended questions for VinDr and MIMIC-CXR, respectively. This shows the effectiveness of grounding on visual evidence before answering as in our proposal.
Textual Perturbation Test. Table 2 reports the sensitivity of each model to textual language bias, measuring the percentage of changed True Positive answers when either switching an anatomy or a disease. The higher the percentage is, the more robust a model is from textual shortcut bias. Our proposed framework achieves the highest sensitivity measure overall, with the ratio of changed answers drastically increasing by up to 3.89% compared to CRG Wan et al. (2025). Enhancing attention to the queried regions, as in our proposal, helps alleviate the textual and visual shortcut learning. Lastly, fine-tuned open-source models are generally more robust than closed-source models like GPT-4-O, showing that fine-tuning LLMs on our adversarial data remedies their hallucination problems.
Visual Perturbation Test. Table 3 reports the visual sensitive score, with the score defined as the number of changed answers when the original image is overlaid with a healthy localized anatomy. Most multi-modal LLMs in our benchmark shift their answer accordingly around 40-70% of the time in our experiments, indicating the big influence of the area of interest on a model’s output. Our method is the most sensitive to the change in the localized area, increasing the number of visually dependent answers by 1.94% compared to Prompt Highlighter. Zhang et al. (2024b)
Ablation studies. Table 4 reports the effects of each component in our LobA framework: the LLaVA Med backbone, Grounded Text Generation module, and the full framework with self-prompting. Applying Grounded Text Generation improves the model’s performance in some categories, showing the effectiveness of localization-aware training.With the self-prompting framework, the accuracy and textual/visual robustness improve greatly, most evidently increasing 4.45% in TPT score, and 2.73% on MIMIC-CXR dataset Yes/No F1 Score.
Qualitative analysis. Fig 6 showcases our LobA and some other LMM’s response on a few qualitative examples, as well as the visual attention map on the input image. We can see that LobA manages to query the attention weights much more intuitively compared to LLaVA-Med or LISA While LLaVA-Med and LISA’s attention are scattered into many irrelevant areas, in the first example LobA’s main attention weights are allocated around the right lower lung area, and the second example is near the left upper lung. Thanks to the grounding generation module, followed by LobA’s self prompting module to automatically re-weight the attention to the area of interest, the model is able to have better visual reasoning capabilities, leading to more accurate answers compared with other methods.